# GitTaskBench: A Benchmark for Code Agents Solving Real-World Tasks Through Code Repository Leveraging

### 26 Aug 2025 v1

@@
Abstract
Beyond scratch coding, exploiting large-scale code repositories (e.g., GitHub) for practical tasks is vital in real-world software development, yet current benchmarks rarely evaluate code agents in such authentic, workflow-driven scenarios. To bridge this gap, we introduce GitTaskBench, a benchmark designed to systematically assess this capability via 54 realistic tasks across 7 modalities and 7 domains. Each task pairs a relevant repository with an automated, human-curated evaluation harness specifying practical success criteria. Beyond measuring execution and task success, we also propose the alpha-value metric to quantify the economic benefit of agent performance, which integrates task success rates, token cost, and average developer salaries. Experiments across three state-of-the-art agent frameworks with multiple advanced LLMs show that leveraging code repositories for complex task solving remains challenging: even the best-performing system, OpenHands+Claude 3.7, solves only 48.15 % of tasks. Error analysis attributes over half of failures to seemingly mundane yet critical steps like environment setup and dependency resolution, highlighting the need for more robust workflow management and increased timeout preparedness. By releasing GitTaskBench, we aim to drive progress and attention toward repository-aware code reasoning, execution, and deployment—moving agents closer to solving complex, end-to-end real-world tasks. The benchmark and code are open-sourced at https://github.com/QuantaAlpha/GitTaskBench.
@@ Introduction
RepoBench (Liu, Xu, and McAuley 2023) 7778 Code Completion
Swe‑Bench‑Verified (Jimenez et al. 2023) 500 Program Repair
LiveCode (Jain et al. 2024) 584 Programming Competitions
MLAgentBench (Huang et al. 2023) 13 ML Tasks
MLE‑Bench (Chan et al. 2025) 72 Kaggle (ML) Tasks
PaperBench (Starace et al. 2025) 20 Paper Code Replication Tasks
GitTaskBench (Ours) 54 User‑centric, Daily‑life Tasks
@@
Aider GPT-4o 5.56 1.85 10.67 492.67 0.0316
GPT-4.1 11.11 7.41 14.83 734.17 0.0355
Claude 3.5 16.67 12.96 7.48 534.00 0.0304
DeepSeekV3 20.37 16.67 7.51 599.64 0.00269
SWE-Agent GPT-4o 17.58 10.19 275.53 1282.70 0.778
GPT-4.1 38.89 31.48 301.11 2098.33 0.661
o3-mini 25.93 20.37 158.45 215.20 0.175
Claude 3.5 41.67 22.23 455.34 943.30 1.38
Claude 3.7 64.81 42.59 552.79 807.63 1.67
DeepSeekV3 18.52 12.04 412.65 1649.82 0.113
Qwen3-32b 7.41 3.70 1445.97 2405.00 -
Qwen3-32b∗† 16.67 11.11 124.15 559.11 -
Llama3.3-70b 25.83 18.52 397.03 1985.64 -
OpenHands GPT-4o 21.30 14.82 760.53 3990.31 1.94
GPT-4.1 55.56 42.59 465.94 1535.47 0.942
o3-mini 29.63 22.22 2523.53 183637.53 3.58
Claude 3.5 53.70 40.74 2858.00 24929.47 8.95
Claude 3.7 72.22 48.15 9501.25 85033.05 29.8
Gemini-2.5-pro 51.85 35.19 760.88 35173.29 2.18
DeepSeekV3 45.37 26.85 4717.78 31957.67 1.31
Qwen3-8b 1.85 1.85 846.26 2045.00 -
Qwen3-14b 11.11 5.56 339.42 2540.17 -
Qwen3-32b 35.19 25.93 591.02 2097.89 -
Qwen3-32b∗† 44.44 29.63 208.00 8755.35 -
Llama3.3-70b 27.78 20.37 132.69 872.93 -
@@ F. Detailed Cases of Code Agents
Primary Category Secondary Subclasses Key Diagnostic Indicators SWE-Agent OpenHands

### 14 Sep 2025 v2

@@
GitTaskBench: A Benchmark for Code Agents Solving Real-World Tasks
Through Code Repository Leveraging
Ziyi Ni1,2 *Huacan Wang1,* ‡Shuo Zhang3 *Shuo Lu1,2Ziyang He4Wang You5
Zhenheng Tang6Yuntao Du7Bill Sun8Hongzhang Liu9,10Sen Hu10Ronghao Chen10
Bo Li6Xin Li11Chen Hu5,‡Binxing Jiao5Daxin Jiang5,‡Pin Lyu2,‡
1UCAS 2CASIA 3BUPT 4NUS 5StepFun 6HKUST 7SDU 8PINAI 9USYD 10PKU 11USTC
* These authors contributed equally to this work.
$\dagger$ Corresponding authors: wanghuacan17@mails.ucas.ac.cn, djiang@stepfun.com, pin.lv@ia.ac.cn
@@ Introduction
RepoBench (Liu, Xu, and McAuley 2023) 7778 Code Completion
Swe‑Bench‑Verified (Jimenez et al. 2023) 500 Program Repair
LiveCode (Jain et al. 2024) 584 Programming Competitions
MLAgentBench (Huang et al. 2023) 13 ML Tasks
MLE‑Bench (Chan et al. 2025) 72 Kaggle (ML) Tasks
PaperBench (Starace et al. 2025) 20 Paper Code Replication Tasks
[…]
@@
[…]
@@ F. Detailed Cases of Code Agents
[…]